[Hallc_running] [Revised Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Fri Apr 20 14:05:01 EDT 2018

Logentry Text:
<p><strong>Update on day and swing shifts of April 19 and owl shift of April 20 </strong></p><p><strong>CEBAF status</strong>:</p><ul><li>Stable beam during day shift (https://logbooks.jlab.org/entry/3563456). 5.7 hours of available beam time. </li><li>Stable beam during swing shift (https://logbooks.jlab.org/entry/3563598). 6.6 hours of available beam time.</li><li>Beam losses during owl shift (https://logbooks.jlab.org/entry/3563768). Operators only able to get stable beam to all Halls by limiting our current to 35muA. Problem might be related to the temperature sensitivity of the beam loss monitors and the temperature gradient during the evening and night. 5.8 hours of available beam time. </li><li>Beam studies including high-current studies for Hall C are planned for Monday, April 23, from about 8:15 a,m. to 3:30 p.m.. Halls A and C are asked to remain locked up during the beam studies. The beam studies require no particular target type and we can choose no target.&nbsp
 ;</li><li>RF recovery planned for Tuesday, April 24. Hall C will be in restricted access due to the tour for the “SURA Board of Trustees Meeting and Meetings of the Standing Committees” (http://www.sura.org/Meetings/2018Spring/). </li><li>During beam studies or RF recovery, more diagnosis software for the C100 will be installed.</li></ul><p><strong>Hall C status</strong>:</p><ul><li>Hall C is configured for the SIDIS experiment (E12-09-017). The HMS is setup for detecting electrons (inclusive measurement) and the SHMS is setup for detecting charged pions and kaons (semi-inclusive measurment). Details on the experimental setup are available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan#Experimental_setup </li><li>For various theta angles in the HMS, the beam current is reduced significantly to keep the pTRIG1 rate below 500kHz. </li><li>During the beam studies on Monday, Brad will upgrade the DAQ and test the upgraded system. </li></ul><p><s
 trong>Data taking</strong>:</p><ul><li>The d!
 etailed plan is available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan </li><li>We have completed data taking for kinematic groups 1 and 2. </li><li>The data taking of the SIDIS experiment with various configuration changes between the runs requires three persons on shift. We have open slots in the shift schedule (https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-2017-2018) and need help with shifts, in particular with the day shift on Sunday and the owl shift of next week. </li><li>On Saturday, Vladimir Berdnikov will take over as run coordinator.</li></ul><p><strong>Hall C Daily RC meeting</strong>: We will meet today at 1:30 p.m. in the conference room located on the second floor of the counting house. Remote participation will be possible via Blue Jeans (Meeting ID: 543903672). You can join via:</p><ul><li>browser: https://bluejeans.com/543903672</li><li>phone: (888) 240-2560 (US toll free), list of international numbe
 rs: https://bluejeans.com/numbers </li></ul><p>There will be no meetings at the weekend.</p>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3563879
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180420/40927505/attachment.html>

More information about the Hallc_running mailing list